# On the capability of high redshift kSZ measurement with galaxy surveys

### v1

@@ Abstract
The kinematic Sunyaev-Zel’dovich (kSZ) effect has been detected at $z<1$ using various techniques and data sets. The ongoing and upcoming spectroscopic galaxy surveys such as DESI (Dark Energy Spectroscopic Instrument) and PFS (Prime Focus Spectrograph) will push the detection beyond $z=1$, and therefore map the baryon distribution at high redshifts. Such detection can be achieved by both the kSZ stacking and tomography methods. While the two methods are theoretically equivalent, they differ significantly in the probed physics and scales, and required data sets. Taking the combination of PFS and ACT (Atacama Cosmology Telescope) as an example, we build mocks of kSZ and galaxies, quantify the kSZ detection S/N, and compare between the two methods. We segment the PFS galaxies into three redshift bins: $1.00.6<z<$ 1.0, $1.61.0<z<$ 1.6, and $2.41.6<z<$ 2.4. For tomography method, our analysis reveals that the two higher redshift bins exhibit significantly higher S/N ratios, with values of 32 and 28, respectively, compared to the first redshift bin, which yielded an S/N of 8. This is attributed to not only the increasing of electron density with redshifts, but also the larger survey volume and the reduced non-linearity, facilitating velocity reconstruction at higher redshifts. Therefore, the capability of the PFS survey to measure high redshift kSZ effect stands as a substantial advantage over other spectroscopic surveys at lower redshift. The S/N of kSZ stacking largely depends on the number of clusters/groups available from another photometric survey. But in general, its S/N is lower than that of kSZ tomography, largely due to CMB instrument noise and error in cluster redshift. Incorporating next-generation CMB surveys like CMB-S4, characterized by significantly reduced instrument noise and improved angular resolution, is expected to enhance tomographic detection by a factor of ten and stacking detection by fivefold. This future high S/N detection holds the promise of not only providing precise constraints on the overall baryon abundance but also initiating a new insight into baryon distribution.
@@ 1 Introduction
In cosmology, the kinematic Sunyaev Zel’dovich (kSZ) effect stands as a pivotal probe for understanding the large-scale baryon distribution of the universe. The kSZ effect, a secondary anisotropy in the CMB, arises from the inverse Compton scattering of CMB photons off a group of free electrons moving with respect to the CMB rest frame. Unlike the thermal Sunyaev-Zel’dovich (tSZ) effect and X-ray emissions, the kSZ effect linearly depends on electron density and is independent of electron temperature, making it a valuable probe for mapping baryon in less dense regions of the universe, such as filaments and voids. This probe has powerful constraining ability in colder and the context of ”missing baryon” problem in the local universe, where observations [1] suggest a $50\%$ deficit relative to the predictions of Big Bang Nucleosynthesis (BBN) [2]. On the other hand, if the baryon distribution can be constraint by other observations, such as tSZ [3], Fast Radio Bursts [4], the kSZ effect encodes valuable information about the growth rate [5, 6] of the universe’s matter content, leveraging the investigation of the dark energy [7], modified gravity [8, 9, 10] and neutrino mass [11].
@@
Despite its significance, the kSZ signal is notably faint compared to other CMB components, presenting a formidable challenge for accurate detection and measurement. A critical aspect of extracting the kSZ signal hinges on its unique dependence on the peculiar velocity field. The first measurement of the kSZ effect by ref. [12] utilized the pairwise method, which capitalizes on the gravitational interaction between cluster pairs moving towards each other. While this method provides insights into the relative motion between clusters, it overlooks the collective motion that could yield additional information. Refs. [13, 14, 15, 16] further developed this method and gain $1.8\sim 5.4\sigma$ detection. The advantage of the pairwise method is not requiring precise knowledge of the clusters’ peculiar velocities and can utilize enormous number of photo-z groups [17, 18, 19, 20].
@@
Advancing the measurement of the kSZ effect involves reconstructing peculiar velocities of clusters, from spectroscopic redshift galaxy surveys [21, 22, 23, 24]. By stacking the CMB signals at cluster locations, the signal-to-noise ratio (S/N) is enhanced. And weighting them by the line-of-sight (LOS) velocities would effectively isolate the kSZ signal from other unrelated CMB components. This approach is particularly effective in studying baryon distributions around clusters. Refs. [25, 26] investigated the effect of various factors on reconstruction efficiency, such as RSD, photometric redshift errors, satellite galaxy fraction, the adoption of incorrect cosmological models, and the scale of smoothing. In addition, they applied this method to a realistic mock, include light cone effect, survey masks and selection functions. In addition, ref. [27] attempted to reconstruct velocity by Convolutional Neural Network and find performance is slightly improved comparing to the conventional method which uses the continuity equation [28].
@@
Existing kSZ measurements have provided useful information on the baryon distribution and evolution in our universe. Ref. [13] sought to quantify the baryon abundance around central galaxies at redshift $z\sim 0.1$ by estimating the pairwise signal using the combination of Planck data and the a Central Galaxy Catalogue samples from the SDSS-DR7. Further refining this measurement, ref. [31] constrained on the baryon fraction by increasing the aperture filter radius and reported an increment in the baryon fraction to 45-55% of the total baryonic content when the filter radius equals to 10 arcmin, albeit with an increasing noise at larger radii. Ref. [32] extended the detection of baryon fraction to higher redshift (z = 0.4), cross-correlating between a projected density field and the squared CMB temperature map. This work achieved a $4.5\sigma$ confirmation that the baryon fraction aligns with BBN prediction. The study continued by ref. [33], which applied a similar methodology across three unWISE galaxy catalog samples peaking at redshifts of 0.6, 1.1, and 1.5. Their results indicated the free gas fraction for the higher redshift bins is about twice higher than the fiducial value. They attributed the abnormally high kSZ signal amplitude to uncertainties in theoretical modeling. In addition, by measuring the pairwise kSZ signal of clusters in redshift span $0.1\sim 0.9$, ref. [17] constrained the gas fraction $f_{\rm gas}=0.086\pm 0.027$ inside the $R_{500}$. Measuring the pairwise kSZ in Fourier space, ref. [34] constrained the gas fraction of LOWZ catalog from BOSS DR12, $f_{\rm gas}=0.17$, while this result has large error contour and has a degeneracy with the assumption of gas profiles. Ref. [35] extracted the kSZ signal of a group catalog from SDSS DR7 [36] and Planck. They divided the catalog into six mass bins and measure the kSZ flux inside $R_{200}$ as a function of halo mass. They found the relationship of kSZ flux and halo mass is self-similar, inferring a mass-independent baryon fraction. They also found the baryon fraction of all mass bin is consistent with the universal value within the errorbars and indicate all “missing baryons” are found on the halo scales. The discussing of “missing baryon” problem has been last twenty years and hasn’t been settled down due to current low measurement significance. Refs. [23, 37, 24] concentrated more on the baryon distribution inside the groups, thanks to the high resolution of ACT survey. They measured the gas density profile by stacking the kSZ signal at the position of a photometric cluster catalog with varying filter size to extract kSZ signal from CMB maps, weighting with the line-of-sight (LOS) velocity reconstructed by another spectroscopic galaxy catalog. They found the gas profile is more extended than that of dark matter and point out the potential of kSZ effect to measuring the feedback in the galaxy formation model [38].
@@
Detection of kSZ through galaxy-CMB cross-correlation so far is limited to $z<1$. Ongoing surveys such as DESI and Euclid and upcoming surveys such as PFS will push the measurement to $z>1$ and reveal the baryons at high redshifts. The capability of kSZ measurement at high redshift is promising due to higher electron number density and potentially better velocity reconstruction. In this study, we forecast the measurement significance of kSZ using the PFS and the CMB surveys such as ACT and CMB-S4. Because FPS provides accurate positions of galaxies with spectroscopic redshifts, which can facilitate the velocity reconstruction, we focus on stacking and tomography methods. These two methods offer complementary insights, measuring the one-halo and two-halo terms of the kSZ signal, respectively. The PFS extends to redshifts of $z=2.4$, enabling high-redshift measurements. Such high-redshift observations can constrain the abundance and evolution of baryons, providing valuable insight in the ‘missing baryon’ problem. Moreover, one-halo term measurements of the kSZ effect could reveal the baryon fraction in clusters as a function of radius, which is sensitive to baryonic processes such supernova and Active Galactic Nuclei (AGN) feedback, aiding in calibrating hydrodynamic simulations. As ref. [39] demonstrates, simulations with varying galaxy formation models align at lower redshifts but diverge significantly beyond $z=2$. Thus, high-redshift kSZ measurements offer invaluable information for understanding the galaxy formation model.
@@
This paper is organized as follows. In section 2, we present the velocity reconstruction algorithm used in both stacking and tomography methods. In section 3 we introduce the N-body simulation CosmicGrowth and the methods to construct the mock galaxy/cluster catalogs and kSZ/CMB maps. Then we show the forecast of the kSZ measurement significance with two method in section 4. In the end, we conclude in section 5.
@@ 2 Velocity reconstruction method
In the stacking method, we obtain the LOS velocity of each cluster $\hat{v}_{{\rm los},i}$, by interpolating the reconstructed field $\hat{\mathbf{v}}(\mathbf{x})$ at their position $\mathbf{x}^{\rm ph}_{i,{\rm cluster}}$.
@@
The two Wiener filters, $W_{1}(k_{\perp},k_{\parallel})$ and $W_{2}(k)$, used to mitigate the noise from RSD and shot noise, have be demonstrate theoretically to be the optimal choice in appendix C. And comparing to conventional Gaussian filter, the Wiener filters can improve the tomography method when $k>0.1h/{\rm Mpc}$ (appendix D).
@@ 3 Mock generation
The PFS is a cosmology survey that offers unique capabilities in astronomical observations due to its large field of view and impressive multiplexing capability, which enables high sampling rates and coverage of a large volume of space [40]. It is designed to explore a previously uncharted epoch in cosmic history, extend to redshift $z=2.4$. The sky coverage is $1400\ {\rm deg^{2}}$. Notably, the PFS survey can cover a comoving volume that is ten times larger than SDSS and twice as large as BOSS, which allows for more detailed studies of cosmic structures and the dark energy component of the universe. The PFS targets a range of galaxy types using optical and near-infrared wavelengths. Specifically, [O II] emission-line galaxies (ELGs), which are useful tracers for conducting an efficient survey out to high redshifts beyond z=1.
In this study, we investigate the potential for detecting the kSZ effect with the PFS ELG catalog. We divide the whole redshift range covered by PFS into three bins: $1.00.6<z<$ 1.0, $1.61.0<z<$ 1.6 and $2.41.6<z<$ 2.4. The comoving intervals of these three bins are almost equivalent. The parameters for the PFS cosmology survey for these three bins are listed in table 1. In this section, we describe the generation of mock kSZ maps, PFS galaxy catalogs and cluster catalogs from the CosmicGrowth simulation using to forecast the kSZ measurement significance.
@@
Table 1: The detailed information of the three redshift bins. The first column shows the redshift range; the second column shows the mean galaxy number density in this redshift range according to PFS survey; the third column show the comoving volume of the redshift bin; the fourth column shows the snapshot in the CosmicGrowth simulation used to conduct the mock kSZ map, galaxy and cluster catalogs in the redshift bin; the fifth column show the HOD model used to generate the ELGs.
@@ 3.1 CosmicGrowth
The CosmicGrowth simulations [42] are a series of high-resolution N-body cosmological simulations designed primarily to study the acceleration of cosmic expansion and the clustering of dark matter and galaxies. The simulations are evolved with an adaptive parallel P3M -body code, while the force between the particles is softened at small scale. In this work, we choose to use the “WMAP_3072_1200” simulation, whose box length is $1200\ {\rm Mpc}/h$. This simulation adopts a $\Lambda\rm CDM$ cosmological model with the cosmological parameters from WMAP satellite, $\Omega_{b}=0.0445$, $\Omega_{c}=0.2235$, $\Omega_{\Lambda}=0.732$, $h=0.71$, $n_{s}=0.968$ and $\sigma_{8}=0.83$. It employs $3072^{3}$ dark matter particles. The dark matter groups are identified by the Friends-of-Friends algorithm and the linking length is set to be 0.2 times the mean particle separation. A FoF group catalog is constructed for each snapshot, providing group mass, positions and peculiar velocities. Specifically, we use the snapshots 2746 ($z=0.811$), 2181 ($z=1.272$) and 1631 ($z=2.023$) to generate mock kSZ map, galaxy and cluster catalogs in the three redshift bins.

### v2

@@ Abstract
The kinematic Sunyaev-Zel’dovich (kSZ) effect has been detected at $z<1$ using various techniques and data sets. The ongoing and upcoming spectroscopic galaxy surveys such as DESI (Dark Energy Spectroscopic Instrument) and PFS (Prime Focus Spectrograph) will push the detection beyond $z=1$, and therefore map the baryon distribution at high redshifts. Such detection can be achieved by both the kSZ stacking and tomography methods. While the two methods are theoretically equivalent, they differ significantly in the probed physics and scales, and required data sets. Taking the combination of PFS and ACT (Atacama Cosmology Telescope) as an example, we build mocks of kSZ and galaxies, quantify the kSZ detection S/N, and compare between the two methods. We segment the PFS galaxies into three redshift bins: $1.00.6<z<$ 1.0, $1.61.0<z<$ 1.6, and $2.41.6<z<$ 2.4. For tomography method, our analysis reveals that the two higher redshift bins exhibit significantly higher S/N ratios, with values of 32 and 28, respectively, compared to the first redshift bin, which yielded an S/N of 8. This is attributed to not only the increasing of electron density with redshifts, but also the larger survey volume and the reduced non-linearity, facilitating velocity reconstruction at higher redshifts. Therefore, the capability of the PFS survey to measure high redshift kSZ effect stands as a substantial advantage over other spectroscopic surveys at lower redshift. The S/N of kSZ stacking largely depends on the number of galaxy groups available from another photometric survey. But in general, its S/N is lower than that of kSZ tomography, largely due to CMB instrument noise and error in galaxy group redshift. Incorporating next-generation CMB surveys like CMB-S4, characterized by significantly reduced instrument noise and improved angular resolution, is expected to enhance tomographic detection by a factor of ten and stacking detection by fivefold. This future high S/N detection holds the promise of not only providing precise constraints on the overall baryon abundance but also initiating a new insight into baryon distribution.
@@ 1 Introduction
In cosmology, the kinematic Sunyaev Zel’dovich (kSZ) effect stands as a pivotal probe for understanding the large-scale baryon distribution of the universe. The kSZ effect, a secondary anisotropy in the CMB, arises from the inverse Compton scattering of CMB photons off a group of free electrons moving with respect to the CMB rest frame. Unlike the thermal Sunyaev-Zel’dovich (tSZ) effect and X-ray emissions, the kSZ effect linearly depends on electron density and is independent of electron temperature, making it a valuable probe for mapping baryon in less dense regions of the universe, such as filaments and voids. This probe has a powerful constraining ability in colder and the context of ”missing baryon” problem in the local universe, where observations [1] suggest a $50\%$ deficit relative to the predictions of Big Bang Nucleosynthesis (BBN) [2]. On the other hand, if the baryon distribution can be constrained by other observations, such as tSZ [3], Fast Radio Bursts [4], the kSZ effect encodes valuable information about the growth rate [5, 6] of the universe’s matter content, leveraging the investigation of the dark energy [7], modified gravity [8, 9, 10] and neutrino mass [11].
@@
Despite its significance, the kSZ signal is notably faint compared to other CMB components, presenting a formidable challenge for accurate detection and measurement. A critical aspect of extracting the kSZ signal hinges on its unique dependence on the peculiar velocity field. The first measurement of the kSZ effect by ref. [12] utilized the pairwise method, which capitalizes on the gravitational interaction between cluster pairs moving towards each other. While this method provides insights into the relative motion between clusters, it overlooks the collective motion that could yield additional information. Refs. [13, 14, 15, 16] further developed this method and gain $1.8\sim 5.4\sigma$ detection. Ref.  [17] constructed the pairwise velocity from the kSZ signal using machine learning, which can break the degeneracy between pairwise velocity and optical depth, and further constrain dark energy and modified gravity theories. The advantage of the pairwise method is that it does not require precise knowledge of the clusters’ peculiar velocities. In addition, apart from the spectroscopic galaxies, the pairwise method can also utilize enormous number of photo-z galaxies/groups [18, 19, 20, 21], although the large redshift uncertainty causes a large suppression of pairwise kSZ signal especially when $r<100\ {\rm Mpc}/h$ [22].
@@
Advancing the measurement of the kSZ effect involves reconstructing peculiar velocities of clusters, from spectroscopic redshift galaxy surveys [23, 24, 25, 26]. By stacking the CMB signals at cluster locations, the signal-to-noise ratio (S/N) is enhanced. Weighting them by the line-of-sight (LOS) velocities would effectively isolate the kSZ signal from other unrelated CMB components. This approach is particularly effective in studying baryon distributions around clusters. Refs. [27, 28] investigated the effect of various factors on reconstruction efficiency, such as RSD, photometric redshift errors, satellite galaxy fraction, the adoption of incorrect cosmological models, and the scale of smoothing. In addition, they applied this method to a realistic mock, include light cone effect, survey masks and selection functions. In addition, ref. [29] attempted to reconstruct velocity by Convolutional Neural Network and found performance is slightly improved compared to the conventional method which uses the continuity equation [30]. In addition, Ref. [31] found the uncertainty in the velocity reconstruction step, which is often neglected, increases the final uncertainty by about 15% at cluster scale.
@@
Moreover, Ref. [34] proposed a novel method called AFR-kSZ tomography. Angular Redshift Fluctuation (ARF), first proposed by [35], is the statistic of sky maps built by projecting redshifts in a thin shell, sensitive to both underlying density and velocity fields. Ref. [34] resorted to the 6dF and SDSS, and foreground-cleaned CMB from Planck to conduct a baryon census in a wide redshift range $50<z<$ 5. They found AFR-kSZ tomography measurement is sensitive to more than half of all baryon in the Universe, with a joint detection $>10\sigma$.
@@
Existing kSZ measurements have provided useful information on the baryon distribution and evolution in our universe. Ref. [13] sought to quantify the baryon abundance around central galaxies at redshift $z\sim 0.1$ by estimating the pairwise signal using the combination of Planck data and the Central Galaxy Catalogue samples from the SDSS-DR7. Further refining this measurement, ref. [36] constrained the baryon fraction by increasing the aperture filter radius and reported an increment in the baryon fraction to 45-55% of the total baryonic content when the filter radius equals to 10 arcmin, albeit with an increasing noise at larger radii. Ref. [37] extended the detection of baryon fraction to higher redshift (z = 0.4), cross-correlating between a projected density field and the squared CMB temperature map. This work achieved a $4.5\sigma$ confirmation that the baryon fraction aligns with BBN prediction. The study continued by ref. [38], which applied a similar methodology across three unWISE galaxy catalog samples peaking at redshifts of 0.6, 1.1, and 1.5. Their results indicated the free gas fraction for the higher redshift bins is about twice as high than the fiducial value. They attributed the abnormally high kSZ signal amplitude to uncertainties in theoretical modeling. In addition, by measuring the pairwise kSZ signal of clusters in redshift span $0.1\sim 0.9$, ref. [18] constrained the gas fraction $f_{\rm gas}=0.086\pm 0.027$ inside the $R_{500}$. Measuring the pairwise kSZ in Fourier space, ref. [39] constrained the gas fraction of LOWZ catalog from BOSS DR12, $f_{\rm gas}=0.17$, while this result has large error contour and has a degeneracy with the assumption of gas profiles. Ref. [40] extracted the kSZ signal of a group catalog from SDSS DR7 [41] and Planck. They divided the catalog into six mass bins and measured the kSZ flux inside $R_{200}$ as a function of halo mass. They found the relationship of kSZ flux and halo mass is self-similar, inferring a mass-independent baryon fraction. They also found the baryon fraction of all mass bin is consistent with the universal value within the errorbars and indicates all “missing baryons” are found on the halo scales. The discussion of “missing baryon” problem has been last twenty years and hasn’t been settled due to current low measurement significance. Refs. [25, 42, 26] concentrated more on the baryon distribution inside the groups, thanks to the high resolution of ACT survey. They measured the gas density profile by stacking the kSZ signal at the position of a photometric cluster catalog with varying filter sizes to extract kSZ signal from CMB maps, weighting with the line-of-sight (LOS) velocity reconstructed by another spectroscopic galaxy catalog. They found the gas profile is more extended than that of dark matter and pointed out the potential of kSZ effect to measure the feedback in the galaxy formation model [43]. Recently, Ref. [44] measured the kSZ profile around clusters using the ACT and DESI photometric galaxies. They found the gas is much more extended than the dark matter at 40 $\sigma$. More interestingly, their results strongly favor high-feedback models in original Illustris instead of low-feedback models in illustrisTNG simulation.
@@
Detection of kSZ through galaxy-CMB cross-correlation so far is limited to $z<1$. Ongoing surveys such as DESI and Euclid and upcoming surveys such as PFS will push the measurement to $z>1$ and reveal the baryons at high redshifts. The capability of kSZ measurement at high redshift is promising due to higher electron number density and potentially better velocity reconstruction. In this study, we forecast the measurement significance of kSZ using the PFS and the CMB surveys such as ACT and CMB-S4. Because FPS provides accurate positions of galaxies with spectroscopic redshifts, which can facilitate the velocity reconstruction, we focus on stacking and tomography methods. These two methods offer complementary insights, measuring the one-halo and two-halo terms of the kSZ signal, respectively. The PFS extends to redshifts of $z=2.4$, enabling high-redshift measurements. Such high-redshift observations can constrain the abundance and evolution of baryons, providing valuable insight in the ‘missing baryon’ problem. Moreover, one-halo term measurements of the kSZ effect could reveal the baryon fraction in clusters as a function of radius, which is sensitive to baryonic processes such as supernova and Active Galactic Nuclei (AGN) feedback, aiding in calibrating hydrodynamic simulations. As ref. [45] demonstrates, simulations with varying galaxy formation models align at lower redshifts but diverge significantly beyond $z=2$. Thus, high-redshift kSZ measurements offer invaluable information for understanding the galaxy formation model.
@@
This paper is organized as follows. In section 2, we present the velocity reconstruction algorithm used in both stacking and tomography methods. In section 3 we introduce the N-body simulation CosmicGrowth and the methods to construct the mock galaxy and galaxy group catalogs and kSZ/CMB maps. Then we show the forecast of the kSZ measurement significance with two methods in section 4. In the end, we conclude in section 5.
@@ 2 Velocity reconstruction method
In the stacking method, we obtain the LOS velocity of each galaxy group $\hat{v}_{{\rm los},i}$, by interpolating the reconstructed field $\hat{\mathbf{v}}(\mathbf{x})$ at their position $\mathbf{x}^{\rm ph}_{i,{\rm group}}$.
@@
The two Wiener filters, $W_{1}(k_{\perp},k_{\parallel})$ and $W_{2}(k)$, used to mitigate the noise from RSD and shot noise, have been demonstrated theoretically to be the optimal choice in appendix C. And comparing to conventional Gaussian filter, the Wiener filters can improve the tomography method when $k>0.1h/{\rm Mpc}$ (appendix D).
@@ 3 Mock generation
The PFS is a cosmology survey that offers unique capabilities in astronomical observations due to its large field of view and impressive multiplexing capability, which enables high sampling rates and coverage of a large volume of space [46]. It is designed to explore a previously uncharted epoch in cosmic history, extending to redshift $z=2.4$. The sky coverage is $1400\ {\rm deg^{2}}$. Notably, the PFS survey can cover a comoving volume that is ten times larger than SDSS and twice as large as BOSS, which allows for more detailed studies of cosmic structures and the dark energy component of the universe. The PFS targets a range of galaxy types using optical and near-infrared wavelengths. Specifically, [O II] emission-line galaxies (ELGs), which are useful tracers for conducting an efficient survey out to high redshifts beyond z=1.
In this study, we investigate the potential for detecting the kSZ effect with the PFS ELG catalog. We divide the whole redshift range covered by PFS into three bins: $1.00.6<z<$ 1.0, $1.61.0<z<$ 1.6 and $2.41.6<z<$ 2.4. The comoving intervals of these three bins are almost equivalent. The parameters for the PFS cosmology survey for these three bins are listed in table 1. In this section, we describe the generation of mock kSZ maps, PFS galaxy catalogs and galaxy group catalogs from the CosmicGrowth simulation used to forecast the kSZ measurement significance.
@@
Table 1: The detailed information of the three redshift bins. The first column shows the redshift range; the second column shows the mean galaxy number density in this redshift range according to PFS survey; the third column shows the comoving volume of the redshift bin; the fourth column shows the snapshot in the CosmicGrowth simulation used to conduct the mock kSZ map, galaxy and galaxy group catalogs in the redshift bin; the fifth column show the HOD model used to generate the ELGs. We assume the density and velocity baryon is following those of dark matter.
@@ 3.1 CosmicGrowth
The CosmicGrowth simulations [48] are a series of high-resolution N-body cosmological simulations designed primarily to study the acceleration of cosmic expansion and the clustering of dark matter and galaxies. The simulations are evolved with an adaptive parallel P3M -body code, while the force between the particles is softened at small scale. In this work, we choose to use the “WMAP_3072_1200” simulation, whose box length is $1200\ {\rm Mpc}/h$. This simulation adopts a $\Lambda\rm CDM$ cosmological model with the cosmological parameters from WMAP satellite, $\Omega_{b}=0.0445$, $\Omega_{c}=0.2235$, $\Omega_{\Lambda}=0.732$, $h=0.71$, $n_{s}=0.968$ and $\sigma_{8}=0.83$. It employs $3072^{3}$ dark matter particles. The dark matter groups are identified by the Friends-of-Friends algorithm and the linking length is set to be 0.2 times the mean particle separation. A FoF group catalog is constructed for each snapshot, providing group mass, positions and peculiar velocities. Specifically, we use the snapshots 2746 ($z=0.811$), 2181 ($z=1.272$) and 1631 ($z=2.023$) to generate mock kSZ map, galaxy and galaxy group catalogs in the three redshift bins.